QueenMidnight16:

assignment question 4 (multiple choice) Two quantities are related, as shown in the table below: x y 2 3 4 4 6 5 8 6 Which equation best represents the relationship? y = 1 over 2x + 2 y = 1 over 2x + 1 y = x + 2 y = 2x + 1

MusicGeek:

For a linear relationship, y=mx+b, m=(y2-y1)/(x2-x1) is a constant, all lines have a constant velocity, called the slope, m. (4-3)/(4-2)=(5-4)/(6-4)=1/2=m. So the slope is 1/2. Then y=x/2+b, using any point we can solve for "b", the y-intercept (the value of y when x=0), I'll use point (2,3). So then you do 3=2/2+b. Then that means that 3=1+b. Which gives you 2=b. So the line is: y=1/2x+2

QueenMidnight16:

thank you

MusicGeek:

you're welcome